Calculate the volume of 0.0500 M EDTA needed to titrate the Ca and Mg in a 0.1612-g sample that is 92.5% dolomite, CaCO 3 ·MgCO 3

(184.4 g/mol)


1
Expert's answer
2022-01-13T08:39:04-0500

Solution:

Calculate the mass of CaCO3 ·MgCO3 in a sample:

Mass of CaCO3 ·MgCO3 = Mass of a sample × %CaCO3 ·MgCO3

Mass of CaCO3 ·MgCO3 = (0.1612 g × 0.925) = 0.14911 g


Calculate the moles of CaCO3 ·MgCO3:

The molar mass of CaCO3 ·MgCO3 is 184.4 g/mol

Hence,

Moles of CaCO3 ·MgCO3 = (0.14911 g) × (1 mol / 184.4 g) = 0.0008086 mol


CaCO3 ·MgCO3 → Ca2+ + Mg2+ + CO32-

According to stoichiometry:

1 mol of CaCO3 ·MgCO3 contain 1 mol of Ca2+ and 1 mol of Mg2+

Hence,

0.0008086 mol of CaCO3 ·MgCO3 contain 0.0008086 mol of Ca2+ and 0.0008086 mol of Mg2+


When calcium ions (Ca2+) are titrated with EDTA (H2Y2-) a relatively stable calcium complex is formed:

Ca2+ + H2Y2- = CaY2- + 2H+

According to stoichiometry:

1 mol of Ca2+ reacts with 1 mol EDTA (H2Y2-)

Hence,

0.0008086 mol of Ca2+ reacts with 0.0008086 mol EDTA (H2Y2-)

Moles 1 of EDTA = 0.0008086 mol


When calcium ions (Mg2+) are titrated with EDTA (H2Y2-) a relatively stable magnesium complex is formed:

Mg2+ + H2Y2- = MgY2- + 2H+

According to stoichiometry:

1 mol of Mg2+ reacts with 1 mol EDTA (H2Y2-)

Hence,

0.0008086 mol of Mg2+ reacts with 0.0008086 mol EDTA (H2Y2-)

Moles 2 of EDTA = 0.0008086 mol


Calculate the moles of EDTA needed to titrate the Ca2+ and Mg2+:

Moles of EDTA = Moles 1 of EDTA + Moles 2 of EDTA

Moles of EDTA = 0.0008086 mol + 0.0008086 mol = 0.0016172 mol


Calculate the volume of EDTA solution:

Volume of EDTA = Moles of EDTA / Molarity of EDTA

Volume of EDTA = (0.0016172 mol) / (0.0500 M) = 0.032344 L = 32.34 mL

Volume of EDTA = 32.34 mL


Answer: The volume of 0.0500 M EDTA needed to titrate the Ca2+ and Mg2+ is 32.34 mL